A recent graduate's student loans total $14,000. If these loans are at 5%, compounded quarterly, for 6 years, what are the quarterly payments? (Round your answer to the nearest cent.)
Answer #1
Quarterly payment = Loan Amount / Present value of annuity of 1
= $ 14,000.00 / 20.62423
= $       678.81
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.0125)^-24)/0.0125 i = 5%/4 = 0.0125
= 20.6242345 n = 6*4 = 24